A schoolteacher is concerned that her students watch more TV than the average American child. She reads that according to the American Academy of Pediatrics (AAP), the average American child watches 4 hours of TV per day
(μ = 4.0 hours).
She records the number of hours of TV each of her six students watch per day. The times (in hours) are 5.5, 4.4, 2.8, 2.5, 4.0, and 4.8.
(a) Test the hypothesis that her students watch more TV than the average American child using a 0.05 level of significance and a one-independent sample t-test. State the value of the test statistic. (Round your answer to three decimal places.)

Answer by stanbon(75887)

the average American child watches 4 hours of TV per day
(μ = 4.0 hours).
She records the number of hours of TV each of her six students watch per day. The times (in hours) are 5.5, 4.4, 2.8, 2.5, 4.0, and 4.8.
sample mean = (sum of data)/6 = 4
sample standard deviation: 1.16
----------------------------
(a) Test the hypothesis that her students watch more TV than the average American child using a 0.05 level of significance and a one-independent sample t-test. State the value of the test statistic. (Round your answer to three decimal places.)
Ho: u = 4
Ha: u > 4 (claim)
-------
t(4) = (4-4)/(1.6/sqrt(6)) = 0
----
p-valu = P(t > 0 when df = 5) = 0.50
Since the p-value is greater than 5%, fail to reject Ho.
Conclusion: The test results do not support the claim.
=======================================================
Cheers,
Stan H.
===============